OpenStudy (anonymous):

x2 –x–6? The 2 after the X is an exponent. This is factoring polynomials( I don't even know what that means) . But I want to know the step by step to solve this type of problem.

OpenStudy (zzr0ck3r):

write exponents like this x^2-x-6

OpenStudy (zzr0ck3r):

\[x^{2}=\]x^2

OpenStudy (zzr0ck3r):

there are a few methods, what are you doing in class?

OpenStudy (anonymous):

I'm in summer, i'm studying to take my SAT.

OpenStudy (anonymous):

i tried to factor it like factoring a quadratic equation but it didn't work.

OpenStudy (zzr0ck3r):

a=1 c=-6 ac=-6 we want two number that multiply and get -6 but add to get b=-1 2*-3=-6 2+-3=-1 so x^2-x-6 = x^2-3x+2x-6 = x(x-3)+2(x-3) = (x-3)(x+2)

OpenStudy (anonymous):

thta's right

OpenStudy (anonymous):

\[x^2 –x–6 =x^2 –3x+2x–6 =x(x-3)+2(x-3) =(x-3)(x+2)\]